# Numerical Homogenization by Continuous Super-Resolution

### v1

@@ Abstract
Finite element methods typically require a high resolution to satisfactorily approximate micro and even macro patterns of an underlying physical model. This issue can be circumvented by appropriate numerical homogenization or multiscale strategies that are able to obtain reasonable approximations on under-resolved scales. In this paper, we study the implicit neural representation and propose a continuous super-resolution network as a numerical homogenization strategy. It can take coarse finite element data to learn both in-distribution and out-of-distribution high-resolution finite element predictions. Our highlight is the design of a local implicit transformer, which is able to learn multiscale features. We also propose Gabor wavelet-based coordinate encodings which can overcome the bias of neural networks learning low-frequency features. Finally, perception is often preferred over distortion so scientists can recognize the visual pattern for further investigation. However, implicit neural representation is known for its lack of local pattern supervision. We propose to use stochastic cosine similarities to compare the local feature differences between prediction and ground truth. It shows better performance on structural alignments. Our experiments show that our proposed strategy achieves superior performance as an in-distribution and out-of-distribution super-resolution strategy.
Index Terms:
numerical homogenization, super-resolution, deep neural networks, finite elements
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
This study paves a new direction of using deep learning-based continuous super-resolution as a numerical homogenization tool. To this end, we address the issue of efficiently simulating physical phenomena characterized by effects on several length scales. Examples of such phenomena are flow and transport through porous media. To reliably predict such phenomena, information on the micro-scale is necessary. However, if, e.g., flow through soil is considered, we are usually only interested in some macroscopic outcome, such as the final macroscopic location of some contaminant in a field. In mathematical terms, such processes are described by partial differential equations (PDEs). More precisely, the equilibrium state of a flow process is characterized by a function $u$ that solves the PDE
@@
$-\operatorname{div}(A\nabla u)&=f&&\quad\text{ in }\Omega,\\ u&=0&&\quad\text{ on }\partial\Omega,\end{aligned}\quad\right.$ (1)
@@
where $\Omega\subset\mathbb{R}^{d}$ is a bounded convex domain in $d$ dimensions, $A\colon\Omega\to\mathbb{R}$ is a scalar-valued coefficient, and $f\colon\Omega\to\mathbb{R}$ is a source or sink. Here, the multi-scale character is encoded in the structure of $A$, which is strongly heterogeneous on a fine scale.
@@
Since problems of the type (1) are not only relevant in soil science [1, 2], oil recovery [3] (and the references therein), drug development [4], and medicine (where the porous medium could, e.g., be a brain [5] or the skin of an individual [6]), but in many disciplines of science and engineering, many approaches in the literature try to tackle them.
@@
Mathematical approaches. From a mathematical perspective, (1) cannot be approximated by standard numerical schemes such as finite element, finite volume, or finite difference schemes on the coarse scale of interest since their respective meshes would need to resolve the fine-scale structure. This would result in an unfeasible number of degrees of freedom and computational demands that forbid the application of these techniques to technically relevant problems. However, there are mathematically rigorous and efficient approaches to tackle this issue, such as [7, 8, 9, 10]. These approaches are based on analytical homogenization theory, see [11, 12], and the references therein, that require the coefficient to exhibit specific properties (such as periodicity or scale separation) that are rarely met in practice. Alternatively, heuristic approaches try to infer easy-to-compute homogeneous problems from statistical considerations [13].
@@
A promising and reliable alternative to the approaches above are computational multi-scale methods, referred to as numerical homogenization. The idea is to generate optimally adapted finite element solutions from a coarse-scale finite element solution enriched by suitable fine-scale information. In particular, such approaches provably work for a vast class of heterogeneous media with minimal structural assumptions, see, e.g., [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Although these methods are highly efficient for solving a multi-scale problem once the enrichment functions are computed, they need substantial pre-computations per choice of $A$.
@@
Our approach. We propose a different approach in the spirit of numerical homogenization. The goal is to learn a mapping from a (potentially incorrect) coarse-scale solution to an improved (upscaled) solution that appropriately represents fine-scale features for a large class of coefficients $A$ and not for each $A$ individually as in the above techniques. That is, we connect the PDE problem to image super-resolution, which is a fundamental task in image processing to recover fine textures in given images. As tools for ill-posed inverse problem, deep learning-based super-resolution approaches have shown remarkable progress in the past decade and have been widely used in digital entertainment, content creation, and fashion design. Recently, researchers have been interested in applying super-resolution to physical problems, including fine-grained climate prediction [27], remote sensing [28], and coastal simulations [29]. Our contributions can be summarized as follows:
@@
  • We propose a numerical homogenization strategy via continuous super-resolution (NH-CSR), incorporating a coarse (and possibly) inaccurate approximation of (1) and the coefficient map $A$ to achieve arbitrary fine approximations via super-resolution.
@@
  • We also employ the multi-scale implicit image function (MS-IIF) for frequency interpolation. It can faithfully preserve low-frequency information while introducing high-frequency details for better visualization.
@@
Figure 1 shows an exemplary comparison of our approach with local implicit image functions (LIIF) [30] on the finite element data super-resolution. We apply the NH-CSR to the low-resolution data and produce multi-scale super-resolved images. Since the model is trained for the setting with an upsampling factor of 16, any upsampling factors that are larger than 16 prescribe an out-of-distribution super-resolution. In the right upper corner (left), we also show the coefficients that are used for the data generation. The remaining plots show visualizations of the upsampled solutions, which are normalized to $[0,255]$, then multiplied by a factor 64 and cut again to the range $[0,255]$ by taking the values modulo 255. These are then used for a coloring using the RGB color space. This is done to highlight pattern differences more clearly. In particular, this leads to the oscillatory structure of the pictures. Overall, we observe that our approach outperforms LIIF in all upsampling scenarios.
The manuscript is structured as follows: In the next section, we discuss the available mathematical tools to tackle (1). Afterward, we mention related works in the context of super-resolution. Section III details our approach regarding the possible application of super-resolution to achieve similar results as numerical homogenization approaches. Finally, Section IV comprises illustrations of our numerical results.
Figure 1: Continuous super-resolution for finite element data. We use the proposed method to apply multi-scale upscaling to the low-resolution image, both in-distribution (orange region, upscaling factor not larger than 16) and out-of-distribution (green region, upscaling factor larger than 16). We see from the comparison that ours achieves better visualization than the state-of-the-art LIIF [30] approach. Note that the actual values are first normalized to [0, 255], then multiplied by a factor of 64, and taken modulo 255 to again be in the value range between 0 and 255. Finally, the result is converted to the RGB color space using the Matplotlib colormap to better visualize differences. This procedure aims at highlighting fine differences in simulation results and leads to the observed wavy patterns.
@@ II-A Mathematical approaches
The basis of numerical homogenization strategies are traditional discretization methods, such as the finite element method. We briefly state the definition of the method and then review some multi-scale methods based on it.
@@
for all appropriate test functions $v$. The finite element method restricts possible functions to approximate $u$ (trial functions) and functions $v$ (test functions) to a finite-dimensional subspace that is built from a decomposition of the domain into a grid $\mathcal{T}_{h}$ consisting of small boxes with side length $h>0$. More precisely, let
$V_{h}:=\left\{\begin{aligned} &v\colon\Omega\to\mathbb{R}\;\colon\;v|_{K}\text{ is a polynomial of}\\ &\text{partial degree }\leq 1\text{ for all }K\in\mathcal{T}_{h}\end{aligned}\right\}\cap C^{0}(\Omega),$
@@
As mentioned above, the finite element method only provides reasonable approximations if $h<\varepsilon$, where $\varepsilon$ is the most miniature scale on which the coefficient varies. However, resolving all details can be a computationally demanding task or even not possible at all for large systems of equations. Therefore, alternative strategies are necessary.
Multi-scale strategies. An alternative to globally resolving fine scales are homogenization-type approaches. Suppose $A$ has a specific structure (e.g., indexed by an explicit, fine parameter $\varepsilon$, periodic, etc.), then homogenization theory states the existence of a homogenized coefficient $A_{0}$ that replaces $A$ in (1) and leads to a solution that presents the macroscopic behavior of the exact solution $u$, see [11, 12]. In particular, a standard finite element method (as discussed above) may now be applied since problematic oscillations have been removed. There are constructive approaches to (approximately) obtain the coefficient $A_{0}$, such as, e.g., the heterogeneous multiscale method, see [9, 10].
Other possibilities are computational strategies that evolve around modifying the space $V_{h}$ to include information about under-resolved scales. Corresponding strategies are, e.g., the multiscale finite element method [7], generalized multiscale finite element methods [14, 15], multiscale spectral generalized finite element methods [16, 17, 18], the localized orthogonal decomposition strategy [19, 20, 21], or gamblets [22, 23]. They often pose fewer assumptions on the structure of the coefficient $A$. For a review article on computational strategies, see [26].
@@ II-B Super-resolution
Implicit neural representation. Implicit neural representation (INR) uses multi-layer perceptrons (MLP) to learn pixels or other signals from coordinates. It has been widely used in 3D shape modeling [31, 32], surface reconstruction [33, 34], novel view rendering [35, 36, 37], etc. For instance, Mildenhall et al. [35] propose to map the camera poses and point clouds to the pixel values via a multi-layer MLP network. They use multiple-view images to optimize the network for implicit feature representation. Instead of using a voxel or point cloud, the implicit neural representation can 1) capture the fine details of scenes for photo-realistic reconstruction and 2) also emit complex 3D representation as a small number of differentiable network parameters.
@@
INR can be regarded as an efficient, continuous, differentiable signal representation. Siren [38] is one of the pioneering approaches that use periodic functions to build sinusoidal representation networks, such that it can solve complex boundary value problems. However, it suffers from signal noise and parameter variations. Therefore, a wavelet implicit neural representation (WIRE) [39] is proposed to utilize Gabor wavelet activation functions to optimize time-frequency compactness. Its results on inverse problems of images and videos show that Gabor wavelets are a universally superior choice for nonlinearities. In the meantime, band-limited coordinate networks (BACON) [40] propose a similar solution as an analytical Fourier spectrum for multi-scale signal representation. It initializes the frequencies of the sinusoidal layers of a multiplicative filter network (MFN) [41] within a limited bandwidth. Multiple layers of outputs are summed together to produce the overall output so that the network learns to fit the signal in a band-limited fashion.
@@
In 3D processing, researchers propose to process 3D data by MLPs as unsigned distance [31, 32], signed distance [33], or occupancy field [34]. For view rendering, NeRF [35] and Gaussian splitting [42] also show great performance in real-time using implicit radiance field rendering. These advances in INR also inspire INR-based modeling of ordinary or partial differential equations. For example, Fourier neural operators [43] parameterize the integral kernel directly in Fourier space with simple multiplications, which leads to a significant reduction of computational costs. Spherical Fourier neural operators [44] generalize the Fourier transform, achieving stable autoregressive long-term forecasting.
@@
Continuous image super-resolution. Image and video super-resolution are ill-posed problems: One low-resolution (LR) image can lead to multiple plausible sharp and clean high-resolution (HR) images. The typical solution is to use paired LR-HR data to optimize restoration models via pixel-based loss functions. Since the seminal super-resolution work of Dong et al. [45], many image super-resolution approaches [46, 47, 48, 49, 50, 51, 52, 53, 54] have adopted an end-to-end neural network to learn the regression model for reconstruction. However, the existing state-of-the-art super-resolution methods can only handle fixed-scale upscaling. For out-of-distribution upscaling, they must be retrained or fine-tuned for optimal results. Continuous super-resolution, on the other hand, combines the recent developments of INR and super-resolution to come up with a grid-based latent space super-resolution. The local implicit image function (LIIF) [30] is one of the pioneering strategies for continuous feature interpolation. That is, a grid-aware MLP network takes both pixel coordinates and global image features for feature expansion and then projects the super-resolved features to the actual pixel values. Such an operation can be further extended to video super-resolution. VideoINR [55] applies INR-based super-resolution to both space and time. Local texture estimators (LTE) [56] improve them by estimating the dominant frequencies to explore the 2D Fourier space feature representation. Local implicit transformers (LITs) [57] are used to integrate the attention mechanism and frequency encoding technique into the INR function. The cross-scale feature aggregation ensures a rich non-local representation. Given the recent success of diffusion models [51] and attention to super-resolution [50], combining INR and diffusion denoising models [58] shows superior performance. CiaoSR [59] proposes a multi-scale attention-in-attention network for continuous super-resolution. Another direction is to use normalizing flow processes [60]. The idea is to learn how to model the ill-posed problem as a super-resolution distribution. LINF [61] learns a coordinate conditional normalizing flow model that can extract the local texture patches for global reconstructions.

### v2

# Multiscale Corrections by Continuous Super-Resolution

@@ Abstract
Finite element methods typically require a high resolution to satisfactorily approximate micro and even macro patterns of an underlying physical model. This issue can be circumvented by appropriate multiscale strategies that are able to obtain reasonable approximations on under-resolved scales. In this paper, we study the implicit neural representation and propose a continuous super-resolution network as a correction strategy for multiscale effects. It can take coarse finite element data to learn both in-distribution and out-of-distribution high-resolution finite element predictions. Our highlight is the design of a local implicit transformer, which is able to learn multiscale features. We also propose Gabor wavelet-based coordinate encodings which can overcome the bias of neural networks learning low-frequency features. Finally, perception is often preferred over distortion so scientists can recognize the visual pattern for further investigation. However, implicit neural representation is known for its lack of local pattern supervision. We propose to use stochastic cosine similarities to compare the local feature differences between prediction and ground truth. It shows better performance on structural alignments. Our experiments show that our proposed strategy achieves superior performance as an in-distribution and out-of-distribution super-resolution strategy.
publicationid: pubid:
@@ I Introduction
This study paves a new direction of using deep learning-based continuous super-resolution as a numerical homogenization tool. To this end, we address the issue of efficiently simulating physical phenomena characterized by effects on several length scales. In this context, the term numerical homogenization refers to a general mathematical strategy that allows to employ fine-scale features of a solution (sometimes referred to as oscillations) in a coarse-scale simulations. Thus, numerical homogenization can in some sense be translated to multiscale correction in what follows. A comparison between classical mathematical homogenization approaches and our super-resolution-based approach is illuminated in the below paragraphs.
We denote our approach continuous, since as opposed to most super-resolution approaches, which are trained on fixed integer scales (e.g., the super-resolution is conducted by factors of 2, 4, …), we allow for an arbitrary scaling and also allow for out-of-distribution resolutions. That is, we may train our model to downscale the simulation $16$ times, and use it to downscale our numerical results $10$ or $20$ times.
@@
Examples of multiscale phenomena with effects on multiple scales are flow and transport through porous media. To reliably predict such phenomena, information on the micro-scale is necessary. However, if, e.g., flow through soil is considered, we are usually only interested in some macroscopic outcome, such as the final macroscopic location of some contaminant in a field. In mathematical terms, such processes are described by partial differential equations (PDEs). More precisely, the equilibrium state of a flow process is characterized by an unknown function $u$ that solves the PDE
@@
$\left\{\begin{aligned} -\operatorname{div}(A\nabla u)&=f&&\quad\text{ in }\Omega,\\ u&=0&&\quad\text{ on }\partial\Omega,\end{aligned}\quad\right$. (1)
@@
in a bounded domain $\Omega\subset\mathbb{R}^{d}$. Here, $A\colon\Omega\to\mathbb{R}$ is a scalar-valued coefficient, and $f\colon\Omega\to\mathbb{R}$ is a source or sink. The multiscale character is encoded in the structure of $A$, which is strongly heterogeneous on a fine scale.
Our main assumption in this manuscript is that $\Omega$ is a rectangular domain covered by some (potentially very fine) pixel mesh on which $A$ varies between two discrete values. The restriction of $A$ can be justified in many cases, e.g., if flow through a porous medium, which results in to two distinct diffusivities, and by a measurement device that is only capable of classifying the medium on the pixel mesh.
@@
Since problems of the type (1) are not only relevant in soil science [ZHANG2023105247, Rupp19], oil recovery [RainerOil] (and the references therein), drug development [https://doi.org/10.1002/zamm.201200196], and medicine (where the porous medium could, e.g., be a brain [https://doi.org/10.1002/pamm.200810201] or the skin of an individual [Reis_2015]), but in many disciplines of science and engineering, many approaches in the literature try to tackle them.
@@
Mathematical approaches. From a mathematical perspective, (1) cannot be approximated by standard numerical schemes such as finite element, finite volume, or finite difference schemes on the coarse scale of interest since their respective meshes would need to resolve the fine-scale structure. This would result in an unfeasible number of degrees of freedom and computational demands that forbid the application of these techniques to technically relevant problems. However, there are mathematically rigorous and efficient approaches to tackle this issue, such as [HouW97, MatS02, EE03, EE05]. These approaches are based on analytical homogenization theory, see [BlaLB23, Tar09], and the references therein, that require the coefficient to exhibit specific properties (such as periodicity or scale separation) that are rarely met in practice. Alternatively, heuristic approaches try to infer easy-to-compute homogeneous problems from statistical considerations [RayRSK18].
@@
A promising and reliable alternative to the approaches above are computational multiscale methods, referred to as numerical homogenization. The idea is to generate optimally adapted finite element solutions from a coarse-scale finite element solution enriched by suitable fine-scale information. In particular, such approaches provably work for a vast class of heterogeneous media with minimal structural assumptions, see, e.g., [EfeGH13, ChuEL18b, BabL11, MaSD22, MaS22, HenP13, MalP14, MalP20, Owh17, OwhS19, Mai21, DonHM23, AltHP21]. Although these methods are highly efficient for solving a multiscale problem once the enrichment functions are computed, they need substantial pre-computations per choice of $A$.
@@
Our approach. We propose a different approach in the spirit of numerical homogenization. The goal is to learn a mapping from a (potentially incorrect) coarse-scale solution to an improved (upscaled) solution that appropriately represents fine-scale features for a large class of coefficients $A$ and not for each $A$ individually as in the above techniques. That is, we connect the PDE problem to image super-resolution, which is a fundamental task in image processing to recover fine textures in given images. As tools for ill-posed inverse problem, deep learning-based super-resolution approaches have shown remarkable progress in the past decade and have been widely used in digital entertainment, content creation, and fashion design. Recently, researchers have been interested in applying super-resolution to physical problems, including fine-grained climate prediction [climate], remote sensing [remote], and coastal simulations [LiuBAR24]. Our contributions can be summarized as follows:
@@
  • We propose a numerical homogenization strategy via continuous super-resolution (NH-CSR), incorporating a coarse (and possibly) inaccurate approximation of (1) and the coefficient map $A$ to achieve reliable approximations via super-resolution.
@@
  • We also employ the multiscale implicit image function (MS-IIF) for frequency interpolation. It can faithfully preserve low-frequency information while introducing high-frequency details for better visualization.
@@
Figure 1 shows an exemplary comparison of our approach with local implicit image functions (LIIF) [LIIF] on the finite element data super-resolution. We apply the NH-CSR to the low-resolution data and produce multiscale super-resolved images. Since the model is trained for the setting with an upsampling factor of 16, any upsampling factors that are larger than 16 prescribe an out-of-distribution super-resolution. In the right upper corner (left), we also show the coefficients that are used for the data generation. The remaining plots show visualizations of the upsampled solutions, which are normalized to $[0,255]$, then multiplied by a factor 64 and cut again to the range $[0,255]$ by taking the values modulo 255. These are then used for a coloring using the RGB color space. This is done to highlight pattern differences more clearly. In particular, this leads to the oscillatory structure of the pictures. Overall, we observe that our approach outperforms LIIF in all upsampling scenarios.
Theoretical grounding. The computational multiscale methods, mentioned above, use the fact that the fine-scale behavior of the solution $u$ can be reconstructed from the coefficientAA. In particular, they correct coarse-scale shape functions in an $A$-dependent fashion to account for the fine-scale behavior. Afterward, they solve a coarse finite element problem with these ‘corrected’ ansatz functions.
Contrasting that approach, we first solve the coarse finite element problem with ‘uncorrected’ shape functions, and correct the solution $u_{h}$ afterward. To that end, our approach receives the coefficient $A$ and the approximated solution $u_{h}$ to generate the fine-scale features. It can, thus, be justified similarly to its mathematical analogs.
In essence, since the PDE of interest is known, only the coefficient $A$ (and right-hand side $f$) is required to obtain a reasonable approximation, see also classical theory on finite element methods, e.g., [Braess, BrennerScott, Bartels, KnabnerA21]. In a similar manner, several works have shown (theoretically and experimentally) that the coefficient-to-solution map can be well-approximated by neural networks, see, e.g., [GeiPRSK21, BhaHKS21, GaoSW21, KutPRS19]. In that sense, our approach even works with more data than theoretically required, as the (easy-to-compute) coarse-scale solution is also taken as input, giving already a very rough approximation of the desired solution as a starting point.
Structure. The manuscript is structured as follows: In the next section, we discuss the available mathematical tools to tackle (1). Afterward, we mention related works in the context of super-resolution. Section III details our approach regarding the possible application of super-resolution to achieve similar results as numerical homogenization approaches. Finally, Section IV comprises illustrations of our numerical results.
Figure 1: Continuous super-resolution for finite element data. We use the proposed method to apply multiscale upscaling to the low-resolution image, both in-distribution (orange region, upscaling factor not larger than 16) and out-of-distribution (green region, upscaling factor larger than 16). We see from the comparison that ours achieves better visualization than the state-of-the-art LIIF [LIIF] approach. Note that the actual values are first normalized to [0, 255], then multiplied by a factor of 64, and taken modulo 255 to again be in the value range between 0 and 255. Finally, the result is converted to the RGB color space using the Matplotlib colormap to better visualize differences. This procedure aims at highlighting fine differences in simulation results and leads to the observed wavy patterns.
@@ II-A Mathematical approaches
The basis of numerical homogenization strategies are traditional discretization methods, such as the finite element method. We briefly state the definition of the method and then review some multiscale methods based on it.
@@
for all appropriate test functions $v$. The finite element method restricts possible functions to approximate $u$ (trial functions) and functions $v$ (test functions) to a finite-dimensional subspace that is built from a decomposition of the domain into a grid $\mathcal{T}_{h}$ consisting of small boxes $K$ with side length $h>0$. More precisely, let
[…]
@@
As mentioned above, the finite element method only provides reasonable approximations if $h<\varepsilon$, where $\varepsilon$ is the most miniature scale on which the coefficient varies. However, resolving all details can be a computationally demanding task or even not possible at all for large systems of equations. Therefore, alternative strategies are necessary.necessary
Multiscale strategies. An alternative to globally resolving fine scales are homogenization-type approaches. Suppose $A$ has a specific structure (e.g., indexed by an explicit, fine parameter $\varepsilon$, periodic, etc.), then homogenization theory states the existence of a homogenized coefficient $A_{0}$ that replaces $A$ in (1) and leads to a solution that presents the macroscopic behavior of the exact solution $u$, see [BlaLB23, Tar09]. In particular, a standard finite element method (as discussed above) may now be applied since problematic oscillations have been removed. There are constructive approaches to (approximately) obtain the coefficient $A_{0}$, such as, e.g., the heterogeneous multiscale method, see [EE03, EE05].
Other possibilities are computational strategies that evolve around modifying the space $V_{h}$ to include information about under-resolved scales. Corresponding strategies are, e.g., the multiscale finite element method [HouW97], generalized multiscale finite element methods [EfeGH13, ChuEL18b], multiscale spectral generalized finite element methods [BabL11, MaSD22, MaS22], the localized orthogonal decomposition strategy [HenP13, MalP14, MalP20], or gamblets [Owh17, OwhS19]. They often pose fewer assumptions on the structure of the coefficient $A$. For a review article on computational strategies, see [AltHP21].
@@ II-B Super-resolution
Implicit neural representation. Implicit neural representation (INR) uses multi-layer perceptrons (MLP) to learn pixels or other signals from coordinates. It has been widely used in 3D shape modeling [shape_1, shape_2], surface reconstruction [surface_1, surface_2], novel view rendering [NERF, structure_1, structure_2], etc. For instance, Mildenhall et al. [NERF] propose to map the camera poses and point clouds to the pixel values via a multi-layer MLP network. They use multiple-view images to optimize the network for implicit feature representation. Instead of using a voxel or point cloud, the implicit neural representation can 1) capture the fine details of scenes for photo-realistic reconstruction and 2) also emit complex 3D representation as a small number of differentiable network parameters.
@@
INR can be regarded as an efficient, continuous, differentiable signal representation. Siren [siren] is one of the pioneering approaches that use periodic functions to build sinusoidal representation networks, such that it can solve complex boundary value problems. However, it suffers from signal noise and parameter variations. Therefore, a wavelet implicit neural representation (WIRE) [wire] is proposed to utilize Gabor wavelet activation functions to optimize time-frequency compactness. Its results on inverse problems of images and videos show that Gabor wavelets are a universally superior choice for nonlinearities. In the meantime, band-limited coordinate networks (BACON) [bacon] propose a similar solution as an analytical Fourier spectrum for multiscale signal representation. It initializes the frequencies of the sinusoidal layers of a multiplicative filter network (MFN) [mfn] within a limited bandwidth. Multiple layers of outputs are summed together to produce the overall output so that the network learns to fit the signal in a band-limited fashion. MIRE [mire] is proposed to use dictionary learning to match the activation of each layer for better signal representation. EVOS [evos] optimize the INR training via the modified evolution algorithm, which reduces $48\%\sim 66\%$ training computation. ViiNeuS [viineus] proposes to use INR on a 3D surface representation. It learns the signed distance field from 2D images for LiDAR datasets.
@@
In 3D processing, researchers propose to process 3D data by MLPs as unsigned distance [shape_1, shape_2], signed distance [surface_1], or occupancy field [surface_2]. For view rendering, NeRF [NERF] and Gaussian splitting [gaussians] also show great performance in real-time using implicit radiance field rendering. These advances in INR also inspire INR-based modeling of ordinary or partial differential equations. For example, Fourier neural operators [fno] parameterize the integral kernel directly in Fourier space with simple multiplications, which leads to a significant reduction of computational costs. Spherical Fourier neural operators [sfno] generalize the Fourier transform, achieving stable autoregressive long-term forecasting.
@@
Continuous image super-resolution. Image and video super-resolution are ill-posed problems: One low-resolution (LR) image can lead to multiple plausible sharp and clean high-resolution (HR) images. The typical solution is to use paired LR-HR data to optimize restoration models via pixel-based loss functions. Since the seminal super-resolution work of Dong et al. [srcnn], many image super-resolution approaches [sr_1, sr_2, sr3, srgan, swinir, stablesr, han, tdan, duf] have adopted an end-to-end neural network to learn the regression model for reconstruction. However, the existing state-of-the-art super-resolution methods can only handle fixed-scale upscaling. For out-of-distribution upscaling, they must be retrained or fine-tuned for optimal results. Continuous super-resolution, on the other hand, combines the recent developments of INR and super-resolution to come up with a grid-based latent space super-resolution. The local implicit image function (LIIF) [LIIF] is one of the pioneering strategies for continuous feature interpolation. That is, a grid-aware MLP network takes both pixel coordinates and global image features for feature expansion and then projects the super-resolved features to the actual pixel values. Such an operation can be further extended to video super-resolution. VideoINR [vinr] applies INR-based super-resolution to both space and time. Local texture estimators (LTE) [lte] improve them by estimating the dominant frequencies to explore the 2D Fourier space feature representation. Local implicit transformers (LITs) [clit] are used to integrate the attention mechanism and frequency encoding technique into the INR function. The cross-scale feature aggregation ensures a rich non-local representation. Given the recent success of diffusion models [stablesr] and attention to super-resolution [swinir], combining INR and diffusion denoising models [diff_inr] shows superior performance. CiaoSR [ciaosr] proposes a multiscale attention-in-attention network for continuous super-resolution. Another direction is to use normalizing flow processes [normalize_flow]. The idea is to learn how to model the ill-posed problem as a super-resolution distribution. LINF [linf] learns a coordinate conditional normalizing flow model that can extract the local texture patches for global reconstructions. Most recently, NeurOp-diff [neuroop-diff] proposes to utilize a diffusion model based neural operator for remote sensing image super-resolution. HIIF [hiif] proposes a hierarchical encoding structure to replace MLP layers for continuous super-resolution. DDIR [ddir] combines the deformation field learning and implicit image function for real-world image super-resolution.
II-C Combining mathematical and super-resolution approaches
Summarizing the above mathematical approaches, we recognize that numerical homogenization methods provably produce accurate solutions under specific model assumptions. However, these methods rely on a computational demanding offline phase for each fine-scale problem. That is, the fine-scale information correcting the coarse-scale solution needs to be calculated individually for each problem.
Machine learning-based super-resolution approaches learn the coarse to fine mapping for a vast class of fine-scale information. However, standard approaches do not respect the fine-scale structure of the individual problem; they rather learn a general fine to coarse mapping that is good for any fine-scale structure.
Our approach bridges this gap between mathematical numerical homogenization and machine learning-based super-resolution: It learns a super-resolution map that contains coarse-to-fine mappings that respect the individual fine-scale information. Thus, our model needs a simple and possibly inaccurate coarse-scale simulation and the fine-scale structure of the problem to generate a fine-scale approximation to the solution. As this approach is very general and efficient, it cannot provide results of the quality of other specifically constructed numerical homogenization tools that invest additional resources for every coefficient separately. Nonetheless, we achieve very good results in a macroscopic sense, which is already very important for multiscale problems. In particular, the results are much better than straight-forward coarse simulations. However, there is (as almost always for machine learning models) no mathematical proof that the result satisfies mathematical accuracy bounds. While one could check the quality of the obtained solution a posteriori, which is a future research prospect, we rather want to provide an intuitive justification for our approach:
The model (1) can be viewed as a description of (slow) fluid flow through soil. The coarse-scale approximation discards fine properties of the flow and is therefore inaccurate. The missing piece is still encoded in the actual soil structure. Since our approach uses both pieces of information (coarse flow and fine-scale structure of porous matrix) and learns to transfer simulation results from a coarse to a fine scale, we can hope to obtain good solutions. This is experimentally justified in Section 4.
Building on this motivation, we further highlight how our method differs from existing super-resolution approaches. Unlike prior works that are limited to fixed-scale super-resolution, our proposed NH-CSR achieves scale-arbitrary super-resolution for both in-distribution and out-of-distribution cases within a single unified model. At its core, NH-CSR employs a multiscale implicit image function, which surpasses existing continuous super-resolution methods in interpolating missing sub-pixels across varying scales. Furthermore, unlike traditional implicit neural representations that require per-signal optimization, NH-CSR adopts a meta-learned, coordinate-based architecture that enables general image representation with a single shared model.